# Port-LLM: A Port Prediction Method for Fluid Antenna based on Large Language Models

### v1

@@ Abstract
The objective of this study is to address the mobility challenges faced by User Equipment (UE) through the implementation of fluid antenna (FA) on the UE side. This approach aims to maintain the time-varying channel in a relatively stable state by strategically relocating the FA to an appropriate port. To the best of our knowledge, this paper introduces, for the first time, the application of large language models (LLMs) in the prediction of FA ports, presenting a novel model termed Port-LLM. The proposed method consists of two primary steps for predicting the moving port of the FA: the first involves utilizing the channel tables that encompass historical channel state information from all movable ports of the FA to forecast the channel tables for subsequent time periods; the second step entails selecting the port of the FA for the forthcoming time based on the predicted channel tables and the known reference channels that require alignment. To enhance the learning capabilities of the LLM model in the context of FA port prediction, we incorporate the Low-Rank Adaptation (LoRA) fine-tuning technology. Furthermore, during the model training phase, we implement the warm-up-aided cosine learning rate (LR) technique to augment the accuracy of the predictions. The simulation results show that our model exhibits strong generalization ability and robustness under different numbers of base station antennas and medium-to-high mobility speeds of UE. In comparison to existing FA port calculation methods, the performance of the port predicted by our model demonstrates superior efficacy. Additionally, our model exhibits lower prediction costs and faster prediction and reasoning speeds.
@@
Fluid antennas, large language models, channel prediction, moving port prediction, Port-LLM.
@@ I Introduction
In recent decades, the implementation of Multiple-Input Multiple-Output (MIMO) technology has significantly enhanced the capacity and reliability of communication systems. Nevertheless, the fixed deployment of current antennas limits the utilization of spatial degrees of freedom (DoF), particularly in environments with little scattering. Additionally, the constraints imposed by the antenna spacing, i.e., the half-wavelength limitation, further restrict the number of antennas that can be deployed within a confined space. This limitation is particularly pronounced on the UE side, where the spatial dimensions are typically small, resulting in inadequate exploitation of narrow space diversity [1, 2, 3]. In contrast, a novel antenna technology known as the fluid antenna (FA) or movable antenna, depending on the hardware implementation, offers the capability to switch among various positions, referred to as “ports”, within a specified area. In this paper, the terms “Fluid antenna” and “Movable antenna” are considered interchangeable. In the subsequent discussion, we will employ the term “Fluid antenna” to encompass both concepts. The position and configuration of the FA can be dynamically adjusted, allowing for greater adaptability [4, 5]. Despite the limited size of the movable area of the FA, the potential for numerous movable ports allows for significant diversity gain across a multitude of spatially dependent ports. Furthermore, the continuous mobility of the FA within a specified area facilitates the optimal exploitation of spatial DoF, thereby enhancing the conditions of the wireless channel.
@@
The issue of mobility, referred to as the curse of mobility [11], has consistently been a significant concern within the field of communications. The movement of the UE can introduce a significant Doppler effect, leading to the obsolescence of the communication channel. If the BS performs precoding design with the outdated CSI, it may result in a decline in system performance. To address this mobility challenge, the work in [11] proposed the Vec Prony method and PAD method for channel prediction. Their findings indicate that it can achieve performance levels comparable to stationary scenarios with unaltered channels. As research in FAs has progressed, the paper [12] proposed utilizing the FA to mitigate mobility issues. Their study introduces an algorithm called “MPMP”, which facilitates the port selection for FA. Empirical results demonstrate that the MPMP method outperforms the Vec Prony algorithm in both medium and high mobility scenarios.
@@
Large Language Models (LLMs), such as GPT-1, GPT-2 [19], GPT-3, and LLaMa, have significantly transformed the fields of natural language processing (NLP) and artificial intelligence (AI). Currently, researchers have initiated investigations into the application of LLMs within the physical layer of wireless communication networks. Notable examples include the Csi-LLM [20] and LLM4CP [21] models, which have been developed for channel prediction, as well as proposals for utilizing LLMs in beam prediction tasks [22]. However, there is a notable gap in the literature regarding the application of LLMs for port prediction in FA. This paper addresses this gap by proposing an FA port prediction model that leverages LLMs. However, given that the extensive datasets utilized for the pre-training of LLMs predominantly consist of diverse textual data, these models lack the capability to interpret wireless communication data. As a result, aligning wireless communication data with natural language data remains a significant challenge when implementing LLMs within the physical layer of wireless communication networks.
@@
In this article, we introduce an FA port prediction model, designated as Port-LLM, to tackle the challenges associated with mobility. Our objective is to maintain a relatively stable channel despite the movement of the UE by relocating the FA to the port predicted by our model at each moment. The foundational architecture of our model is derived from the pre-trained GPT-2 [19] framework, and we employ LoRA [23] technology to fine-tune the loaded GPT-2 model. The research presented in [12] indicates that the critical factor in selecting ports for FA lies in the precise forecasting of CSI pertaining to all movable ports of the FA on the UE side at any given moment. As a result, our proposed LLM-based FA port prediction model utilizes the CSI corresponding to all movable ports on the UE side from the preceding $T$ moments, referred to as the “channel tables,” as input. Subsequently, the proposed Port-LLM is employed to forecast the channel tables for the subsequent $F$ moments. Ultimately, the moving ports of the FA at the subsequent $F$ moments are derived from the channel tables predicted by our model, in conjunction with the known reference channels that require alignment. Generally, the process of utilizing the proposed Port-LLM for FA port prediction can be categorized into two primary phases: the first phase involves channel prediction, while the second phase pertains to port selection.
The main contributions of this paper are as follows:
@@
  • To ensure that the wireless communication data pertinent to this task is aligned with the data patterns of the pre-trained LLM, we developed a specialized module for data processing. Concurrently, we conducted LoRA fine-tuning on the GPT-2 model. By employing low-rank matrix techniques, the LoRA fine-tuning approach significantly diminishes the number of parameters required for retraining our model on this specific task, while preserving the knowledge acquired during the pre-training phase of the GPT-2 model.
  • We introduce the warm-up-aided cosine LR scheduler within the training process of the proposed model, which greatly improves the prediction accuracy and convergence speed of our model.
[…]
@@ II System model
We consider a time division duplexing (TDD) system, where a BS with an $N_{y}\times N_{z}$ uniform planar array (UPA) serves a certain UE that is equipped with an FA. The setup is illustrated in Fig. 1.
@@
$\ \mathbf{r}^{\textrm{rx}}=\left[\begin{array}[]{c}\sin\theta_{\text{EOA}}\cos\phi_{\text{AOA}}\\ \sin\theta_{\text{EOA}}\sin\phi_{\text{AOA}}\\ \cos\theta_{\text{EOA}}\\ \end{array}\right],$ (6)
@@
$\alpha_{p}=\begin{cases}\sqrt{\frac{1}{K_{R}+1}},&\ p\in\text{NLoS},\\ \sqrt{\frac{K_{R}}{K_{R}+1}},&\ p\in\text{LoS},\\ \end{cases}$ (8)
@@
$\left(n_{\text{opt}},m_{\text{opt}}\right)=\text{unravel\_index}\bigg{(}&\text{argmin}\bigg{(}\sum_{i=1}^{N_{t}}\Big{|}\mathbf{S}_{i}\left(t+\varDelta t\right)-\\ &\mathbf{H}_{i}\left(t\right)\Big{|}\bigg{)},\left(N,M\right)\bigg{)}\end{aligned}.$ (19)
@@ III Port-LLM
In this section, we propose a model for predicting the moving port of FA, referred to as Port-LLM, which is grounded in LLMs technology. The primary objective of our model is to ensure that the channel remains relatively stable by relocating the FA on the UE side to the anticipated port while the UE is in motion. To accomplish this, we initially employ the proposed Port-LLM model to forecast the channel tables $\mathbf{\hat{S}}=\left\{\mathbf{\hat{S}}_{1},\cdots,\mathbf{\hat{S}}_{F}\right\}\in\mathbb{C}^{F\times N\times M}$ for the subsequent $F$ moments by utilizing the channel tables $\mathbf{S}=\left\{\mathbf{S}_{1},\cdots,\mathbf{S}_{T}\right\}\in\mathbb{C}^{T\times N\times M}$ from the preceding $T$ moments. Subsequently, we proceed to utilize the predicted channel tables $\mathbf{\hat{S}}\in\mathbb{C}^{F\times N\times M}$ and the known reference channel $\mathbf{H}_{\text{ref}}\in\mathbb{C}^{F\times N\times M}$ to obtain the moving ports $\mathbf{P}=\left\{\mathbf{p}_{1},\cdots,\mathbf{p}_{F}\right\}\in\mathbb{R}^{F\times 2\times 1}$ for the subsequent $F$ moments.
@@ III-A2 Input Embedding
To facilitate the extraction of features for subsequent modeling, we initially employ a linear layer to modify the dimensions of the tensor, thereby transforming $\mathbf{\bar{S}}_{r}\in\mathbb{R}^{T\times N\times M}$ and $\mathbf{\bar{S}}_{i}\in\mathbb{R}^{T\times N\times M}$ into $\mathbf{\tilde{S}}_{r}\in\mathbb{R}^{T\times d_{\text{model}}}$ and $\mathbf{\tilde{S}}_{i}\in\mathbb{R}^{T\times d_{\text{model}}}$, respectively. $d_{\text{model}}$ is the feature dimension of the column vector in the matrix $\mathbf{\tilde{S}}_{r}$ or $\mathbf{\tilde{S}}_{i}$. Following this transformation, we implement the multi-head attention module [25] to extract features from both the real and imaginary components along the temporal dimension. For illustrative purposes, we will focus on the processing of the real component data $\mathbf{\bar{S}}_{r}$ as a example. In this context, for each head $k\in\left\{1,\cdots,K\right\}$ within the module, we define the query matrix, key matrix, and value matrix as $\mathbf{Q}_{k}^{r}=\mathbf{\tilde{S}}_{r}\mathbf{W}_{k}^{Q}$, $\mathbf{K}_{k}^{r}=\mathbf{\tilde{S}}_{r}\mathbf{W}_{k}^{K}$ and $\mathbf{V}_{k}^{r}=\mathbf{\tilde{S}}_{r}\mathbf{W}_{k}^{V}$, respectively. The reprogramming operation in each attention head is defined as
@@ III-A3 Backbone network
The LLM selected for our study is the GPT-2 model [19]. We implement LoRA fine-tuning on the pre-trained GPT-2 model. LoRA fine-tuning is based on the intrinsic low-rank characteristics of the LLMs. It simulates full parameter fine-tuning by adding bypass matrices, aiming to achieve lightweight fine-tuning [23]. In particular, we exclusively conduct LoRA fine-tuning and retraining on the $\mathbf{Q}$ and $\mathbf{V}$ computations within the multi-head attention component of the GPT-2 model, while keeping the remaining model parameters frozen. This approach can substantially reduce the computational resources necessary for our model retraining and better utilize the knowledge acquired by GPT-2 during its pre-training phase. It is assumed that the input data for the module requiring LoRA fine-tuning is denoted by $\mathbf{Z}$. The process of LoRA fine-tuning is outlined as follows:
$\mathbf{Q}=\mathbf{W}_{Q}\mathbf{Z}+\mathbf{B}_{Q}\mathbf{A}_{Q}\mathbf{Z}+\mathbf{b}_{Q},$ (22)
$\mathbf{V}=\mathbf{W}_{V}\mathbf{Z}+\mathbf{B}_{V}\mathbf{A}_{V}\mathbf{Z}+\mathbf{b}_{V},$ (23)
where $\mathbf{W}_{Q},\mathbf{W}_{V}\in\mathbb{R}^{d_{m}\times d}$ are the weights of the pre-trained GPT-2 model, which remain constant and are not subject to gradient updates throughout the training process. $\mathbf{b}_{Q}$ and $\mathbf{b}_{V}$ are the biases of the loaded model, which also remain fixed. $\mathbf{B}_{Q},\mathbf{B}_{V}\in\mathbb{R}^{d_{m}\times r}$ and $\mathbf{A}_{Q},\mathbf{A}_{V}\in\mathbb{R}^{r\times d}$ are learnable parameters. Furthermore, $r\ll\min\left(d_{m},d\right)$, resulting in negligible additional inference delay during model prediction when employing LoRA fine-tuning. As illustrated in Fig. 2, we employ random Gaussian initialization for parameter $\mathbf{A}_{Q}$ and $\mathbf{A}_{V}$, and zero initialization for parameter $\mathbf{B}_{Q}$ and $\mathbf{B}_{V}$ prior to the commencement of our model training.
Generally, the data $\mathbf{X}\in\mathbb{R}^{T\times 2\times d_{\text{model}}}$ is processed through a linear layer to modify its dimensionality to $\mathbf{\tilde{X}}\in\mathbb{R}^{F\times d_{\text{model}}}$ prior to being fed into the backbone network. Subsequently, the data is integrated into the backbone network, where the following procedure occurs:

### v2

@@ Abstract
The objective of this study is to address the mobility challenges faced by user equipment (UE) through the implementation of fluid antenna (FA) on the UE side. This approach aims to maintain the time-varying channel in a relatively stable state by strategically relocating the FA to an appropriate port. To the best of our knowledge, this paper introduces, for the first time, the application of large language models (LLMs) in the prediction of FA ports, presenting a novel model termed Port-LLM. Our proposed method for predicting the moving port of the FA is a two-step prediction method. To enhance the learning efficacy of our proposed Port-LLM model, we integrate low-rank adaptation (LoRA) fine-tuning technology. Additionally, to further exploit the natural language processing capabilities of pre-trained LLMs, we propose a framework named Prompt-Port-LLM, which is constructed upon the Port-LLM architecture and incorporates prompt fine-tuning techniques along with a specialized prompt encoder module. The simulation results show that our proposed models all exhibit strong generalization ability and robustness under different numbers of base station antennas and medium-to-high mobility speeds of UE. In comparison to existing methods, the performance of the port predicted by our models demonstrates superior efficacy. Moreover, both of our proposed models achieve millimeter-level inference speed.
@@
Fluid antennas, large language models, channel prediction, moving port prediction, Port-LLM, Prompt-Port-LLM.
@@ I Introduction
In recent decades, the implementation of multiple-input multiple-output (MIMO) technology has significantly enhanced the capacity and reliability of communication systems. Nevertheless, the fixed deployment of current antennas limits the utilization of spatial degrees of freedom (DoF), particularly in environments with little scattering. Additionally, the constraints imposed by the antenna spacing, i.e., the half-wavelength limitation, further restrict the number of antennas that can be deployed within a confined space. This limitation is particularly pronounced on the user equipment (UE) side, where the spatial dimensions are typically small, resulting in inadequate exploitation of narrow space diversity [1, 2, 3]. In contrast, a novel antenna technology known as the fluid antenna (FA) or movable antenna, depending on the hardware implementation, offers the capability to switch among various positions, referred to as “ports”, within a specified area. The position and configuration of the FA can be dynamically adjusted, allowing for greater adaptability [4, 5, 6, 7]. Despite the limited size of the movable area of the FA, the potential for numerous movable ports allows for significant diversity gain across a multitude of spatially dependent ports. Furthermore, the continuous mobility of the FA within a specified area facilitates the optimal exploitation of spatial DoF, thereby enhancing the conditions of the wireless channel.
@@
The issue of mobility, referred to as the curse of mobility [13], has consistently been a significant concern within the field of communications. The movement of the UE can introduce a significant Doppler effect, leading to the obsolescence of the communication channel. If the BS performs precoding design with the outdated CSI, it may result in a decline in system performance. To address this mobility challenge, the work in [13] proposed the Vec Prony method and the Prony based angular-delay domain (PAD) method for channel prediction. Their findings indicate that it can achieve performance levels comparable to stationary scenarios with unaltered channels. As research in FAs has progressed, the paper [14] proposed utilizing the FA to mitigate mobility issues. Their study introduces a matrix pencil-based moving port (MPMP) prediction method, which facilitates the port selection for FA. Empirical results demonstrate that the MPMP method outperforms the Vec Prony algorithm in both medium and high mobility scenarios.
@@
Large language models (LLMs), such as GPT-1, GPT-2 [21], GPT-3, and LLaMa, have significantly transformed the fields of natural language processing (NLP) and artificial intelligence (AI). Compared to other neural network-based models, models based on LLMs not only possess excellent NLP capabilities but also have inherent sequence modeling capabilities acquired during the pre-training phase. Currently, researchers have initiated investigations into the application of LLMs within the physical layer of wireless communication networks. Notable examples include the Csi-LLM [22] and LLM4CP [23] models, which have been developed for channel prediction, as well as proposals for utilizing LLMs in beam prediction tasks [24]. However, there is a notable gap in the literature regarding the application of LLMs for port prediction in FA. This paper addresses this gap by proposing an FA port prediction model that leverages LLMs. However, given that the extensive datasets utilized for the pre-training of LLMs predominantly consist of diverse textual data, these models lack the capability to interpret wireless communication data. As a result, aligning wireless communication data with natural language data remains a significant challenge when implementing LLMs within the physical layer of wireless communication networks.
@@
In this article, we introduce an FA port prediction model, designated as Port-LLM, to tackle the challenges associated with mobility. Our objective is to maintain a relatively stable channel despite the movement of the UE by relocating the FA to the port predicted by our model at each moment. The foundational architecture of our model is derived from the pre-trained GPT-2 [21] framework, and we employ low- rank adaptation (LoRA) [25] technology to fine-tune the loaded GPT-2 model. The research presented in [14] indicates that the critical factor in selecting ports for FA lies in the precise forecasting of CSI pertaining to all movable ports of the FA on the UE side at any given moment. As a result, our proposed LLM-based FA port prediction model utilizes the CSI corresponding to all movable ports on the UE side from the preceding $T$ moments, referred to as the “channel tables,” as input. Subsequently, the proposed Port-LLM is employed to forecast the channel tables for the subsequent $F$ moments. Ultimately, the moving ports of the FA at the subsequent $F$ moments are derived from the channel tables predicted by our model, in conjunction with the known reference channels that require alignment. Generally, the process of utilizing the proposed Port-LLM for FA port prediction can be categorized into two primary phases: the first phase involves channel prediction, while the second phase pertains to port selection.
Furthermore, we propose a model termed Prompt-Port-LLM, which is based on prompt fine-tuning. This model distinguishes itself from our previously proposed Port-LLM by employing dynamic prompts for the fine-tuning of the LLMs, whereas the Port-LLM utilizes LoRA fine-tuning for the same purpose. Based on retaining the modeling capabilities of the pre-trained LLMs used by the Port-LLM model, our proposed Prompt-Port-LLM model further explores its potential in NLP. In the proposed Prompt-Port-LLM model, we design specialized dynamic prompts and a specialized prompt encoding module to assist in training. The dynamic prompts can provide real-time task knowledge based on the input data, while the prompt encoding module transforms static prompts into optimizable continuous embeddings, thereby further enhancing the performance of our model. The main contributions of this paper are as follows:
@@
  • To ensure that the wireless communication data pertinent to this task is aligned with the data patterns of the pre-trained LLM, we develop the specialized data processing module, input embedding module and output projection module. Meanwhile, we conduct LoRA fine-tuning on the GPT-2 model. By employing low-rank matrix techniques, the LoRA fine-tuning approach significantly reduces the number of parameters required for retraining our model on this specific task compared to the full fine-tuning technique, while preserving the knowledge acquired during the pre-training phase of the GPT-2 model.
  • In order to further leverage the NLP capabilities of pre-trained LLMs, we also propose a framework termed Prompt-Port-LLM, which is built upon the Port-LLM architecture and incorporates prompt fine-tuning technique. Within the Prompt-Port-LLM framework, we design the specialized dynamic prompts and a prompt encoding module to assist in model training, thereby further enhancing model performance.
Notation: We use boldface to denote matrices and vectors. $\mathbb{R}^{n}$ and $\mathbb{C}^{n}$ denote the spaces of $n$-dimension real and complex numbers, respectively. $\left(\cdot\right)^{T}$ represents the transpose. $\text{argmin}\left(\cdot\right)$ refers to the input parameter that minimizes the objective function. $\text{unravel\_index}\left(p,\left(n,m\right)\right)$ denotes the multi-dimensional coordinates associated with the integer $p$ within an $n\times m$ dimensional matrix. $\left|\cdot\right|$ is the absolute value, $\lVert\cdot\rVert$ represents the Euclidean norm, and $\mathbb{E}\left[\cdot\right]$ represents the expectation operator. $\lceil\cdot\rceil$ is the rounding up operation.
@@ II System model
We consider a time division duplexing (TDD) system, where a BS with an $N_{y}\times N_{z}$ uniform planar array (UPA) serves a certain UE that is equipped with an FA. The UE-side FA can be dynamically reconfigured by mechanical actuation or electronic switching control, allowing its radiating unit to quickly switch operating port positions within its preset movable area. Fig. 1 is the downlink (DL) wireless communication system with FA at the UE side.
@@
[…]
@@
[…]
@@
$\begin{aligned} \left(n_{\text{opt}},m_{\text{opt}}\right)=\text{unravel\_index}\bigg{(}&\text{argmin}\bigg{(}\sum_{i=1}^{N_{t}}\Big{|}\mathbf{S}_{i}\left(t+\varDelta t\right)-\\ &\mathbf{H}_{i}\left(t\right)\Big{|}\bigg{)},\left(N,M\right)\bigg{)}\end{aligned}.$ (19)
@@ III Port-LLM
In this section, we propose a model for predicting the moving port of FA, referred to as Port-LLM, which is grounded in LLMs technology. The primary objective of our model is to ensure that the channel remains relatively stable by relocating the FA on the UE side to the anticipated port while the UE is in motion. In the subsequent model parameters, in order to simplify the notation, we set the number of antennas on the BS side $N_{t}$ to 1. Note that the simulation validations are performed in multi-antenna setting. To accomplish the above objective, we initially employ the proposed Port-LLM model to forecast the channel tables $\mathbf{\hat{S}}=\left\{\mathbf{\hat{S}}_{1},\cdots,\mathbf{\hat{S}}_{F}\right\}\in\mathbb{C}^{F\times N\times M}$ for the subsequent $F$ moments by utilizing the channel tables $\mathbf{S}=\left\{\mathbf{S}_{1},\cdots,\mathbf{S}_{T}\right\}\in\mathbb{C}^{T\times N\times M}$ from the preceding $T$ moments. Subsequently, we proceed to utilize the predicted channel tables $\mathbf{\hat{S}}\in\mathbb{C}^{F\times N\times M}$ and the known reference channel $\mathbf{H}_{\text{ref}}\in\mathbb{C}^{F\times N\times M}$ to obtain the moving ports $\mathbf{P}=\left\{\mathbf{p}_{1},\cdots,\mathbf{p}_{F}\right\}\in\mathbb{R}^{F\times 2\times 1}$ for the subsequent $F$ moments.
@@ III-A Network Architecture
The network architecture of our proposed Port-LLM model primarily comprises the data processing, input embedding, backbone network, and output projection modules.
@@ III-A2 Input Embedding
For feature extraction in subsequent modeling, the real part $\mathbf{\bar{S}}_{r}\in\mathbb{R}^{T\times N\times M}$ and imaginary part $\mathbf{\bar{S}}_{i}\in\mathbb{R}^{T\times N\times M}$ are first processed separately through a shared feature extraction module (referred to as the “Shared Module”). This shared module consists of two key components: an encoder module (“Encoder Block”) and a multi-head attention module [27]. The encoder module is responsible for dimensionality reduction and feature compression of the input data, with its detailed architecture illustrated in Fig. 3. The multi-head attention module, on the other hand, is primarily employed to extract spatio-temporal features from both the real and imaginary components. Given that the real and imaginary data undergo identical processing within the “Shared Module”, we focus on the real component as an illustrative example for clarity in the following discussion.
The processing sequence for the real part data $\mathbf{\bar{S}}_{r}\in\mathbb{R}^{T\times N\times M}$ within the Encoder Block module is delineated as follows: initially, the input data is subjected to a sequential passage through two DownSampling modules. Each of these modules reduces the spatial dimensions (length and width) of the data by half, while maintaining a constant number of channels. The detailed procedure for the DownSampling process is illustrated below:
$\mathbf{\mathring{S}}_{r}$ $=\text{DownSampling}\left(\mathbf{\bar{S}}_{r}\right)$ (20)
$=\text{Conv2d}\left(\mathbf{\bar{S}}_{r}\right)+\text{Conv\#}2\left(\text{Conv\#}1\left(\mathbf{\bar{S}}_{r}\right)\right),$
where $\mathbf{\mathring{S}}_{r}\in\mathbb{R}^{T\times\frac{N}{2}\times\frac{M}{2}}$, and the skip connection in the DownSampling block is introduced to mitigate feature loss. On the right-hand side of the equation,
  • The first convolution operation (Conv2d) performs spatial downsampling, halving the length and width of the input data. Its parameters are: kernel size = 1, stride = 2, padding = 1.
  • Both Conv#1 and Conv#2 operations consist of a Conv2d, followed by BatchNorm2d and LeakyReLU activation. However, they serve distinct purposes: Conv#1 operation performs downsampling (halving spatial dimensions) with parameters: kernel size = 3, stride = 2, padding = 1; Conv#2 operation enhances feature representation while preserving input dimensionality and its parameters are: kernel size = 3, stride = 1, padding = 1.
Following two DownSampling procedures, the dimensionality of the real data is reduced from an initial value of $T\times N\times M$ to $T\times\lceil\frac{N}{4}\rceil\times\lceil\frac{M}{4}\rceil$. Subsequently, the multidimensional feature data is transformed into a one-dimensional vector through a flattening operation (Flatten). This one-dimensional feature vector undergoes the processing via a linear layer (Linear) and a LeakyReLU activation function, ultimately yielding the output $\mathbf{\tilde{S}}_{r}\in\mathbb{R}^{T\times d_{\text{model}}}$. $d_{\text{model}}$ is the feature dimension of the column vector in the matrix $\mathbf{\tilde{S}}_{r}$. Similarly, after the Encoding Block module, the imaginary component is transformed from $\mathbf{\bar{S}}_{i}\in\mathbb{R}^{T\times N\times M}$ to $\mathbf{\tilde{S}}_{i}\in\mathbb{R}^{T\times d_{\text{model}}}$.
In the multi-head attention module, for each head $k\in\left\{1,\cdots,K\right\}$ within the module, taking the real part data $\mathbf{\tilde{S}}_{r}$ as an example, we define the query matrix, key matrix, and value matrix as $\mathbf{Q}_{k}^{r}=\mathbf{\tilde{S}}_{r}\mathbf{W}_{k}^{Q}$, $\mathbf{K}_{k}^{r}=\mathbf{\tilde{S}}_{r}\mathbf{W}_{k}^{K}$ and $\mathbf{V}_{k}^{r}=\mathbf{\tilde{S}}_{r}\mathbf{W}_{k}^{V}$, respectively. The reprogramming operation in each attention head is defined as
@@
Before the data $\mathbf{X}\in\mathbb{R}^{T\times 2\times d_{\text{model}}}$ is input to the backbone network, the data needs to go through the Input Projection module for further feature extraction and dimension transformation. The processing flow of the Input Projection module is shown below:
$\mathbf{\tilde{X}}=\text{rearrange}\left(\text{Linear}\left(\text{GELU}\left(\text{Linear}\left(\text{rearrange}\left(\mathbf{X}\right)\right)\right)\right)\right),$ (23)
where $\mathbf{\tilde{X}}\in\mathbb{R}^{F\times d_{\text{model}}}$. The hidden layer dimension of the two-layer Linear neural network is $d_{l}$. GELU (Gaussian Error Linear Unit) is the nonlinear activation function. $\text{rearrange}\left(\cdot\right)$ function implements the dimension transformation of the feature tensor.
@@ III-A3 Backbone network
The LLM selected for our study is the GPT-2 model [21]. We implement LoRA fine-tuning on the pre-trained GPT-2 model. LoRA fine-tuning is based on the intrinsic low-rank characteristics of the LLMs. It simulates full parameter fine-tuning by adding bypass matrices, aiming to achieve lightweight fine-tuning [25]. In particular, we exclusively conduct LoRA fine-tuning and retraining on the query matrix $\mathbf{Q}_{k}$ and the value matrix $\mathbf{V}_{k}$ computations within the multi-head attention component of the GPT-2 model, while keeping the remaining model parameters frozen. This approach can substantially reduce the computational resources necessary for our model retraining and better utilize the knowledge acquired by GPT-2 during its pre-training phase. It is assumed that the input data for the module requiring LoRA fine-tuning is denoted by $\mathbf{Z}$. The process of LoRA fine-tuning is outlined as follows:
$\mathbf{Q}_{k}=\mathbf{W}_{k}^{Q}\mathbf{Z}+\mathbf{B}_{k}^{Q}\mathbf{A}_{k}^{Q}\mathbf{Z}+\mathbf{b}_{k}^{Q},$ (24)
$\mathbf{V}_{k}=\mathbf{W}_{k}^{V}\mathbf{Z}+\mathbf{B}_{k}^{V}\mathbf{A}_{k}^{V}\mathbf{Z}+\mathbf{b}_{k}^{V},$ (25)
where $\mathbf{W}_{k}^{Q},\mathbf{W}_{k}^{V}$ are the weights of the pre-trained GPT-2 model, which remain constant and are not subject to gradient updates throughout the training process. $\mathbf{b}_{k}^{Q}$ and $\mathbf{b}_{k}^{V}$ are the biases of the loaded model, which also remain fixed. $\mathbf{B}_{k}^{Q},\mathbf{B}_{k}^{V}\in\mathbb{R}^{d_{m}\times r}$ and $\mathbf{A}_{k}^{Q},\mathbf{A}_{k}^{V}\in\mathbb{R}^{r\times d}$ are learnable parameters. Furthermore, $r\ll\min\left(d_{m},d\right)$, resulting in negligible additional inference delay during model prediction when employing LoRA fine-tuning. As illustrated in Fig. 2, we employ random Gaussian initialization for parameter $\mathbf{A}_{k}^{Q}$ and $\mathbf{A}_{k}^{V}$, and zero initialization for parameter $\mathbf{B}_{k}^{Q}$ and $\mathbf{B}_{k}^{V}$ prior to the commencement of our model training.
Generally, the data $\mathbf{\tilde{X}}\in\mathbb{R}^{F\times d_{\text{model}}}$ is integrated into the backbone network, where the following procedure occurs: